(1 point) Randomly selected 20 student cars (population 1) have ages with a mean of 7.8 years and a standard deviation of 3.4 years, while randomly selected 22 faculty cars (population 2) have ages with a mean of 5.2 years and a standard deviation of 3.5 years. (For the purposes of this exercise, the two populations are assumed to be normally distributed.) 1. Use a 0.04 significance level to test the claim that student cars are older than faculty cars. The test statistic is The P-value (using the approximate number of degrees of freedom) is Is there sufficient evidence to support the claim that student cars are older than faculty cars? OA. Yes O B. No 2. Construct a 96% confidence interval estimate of the difference μι approximate value for the number of degrees of freedom.) μ2 where 11 s the mean age of student cars and μ2 is the mean age of fa cult cars set e
Answer #1

1. H0: The student cards are not older than faculty cars
H1: The student cards are older than faculty cars
Let the los be alpha = 4%
Pooled SD 19*3.42 +21*3.5 20+22-2 = 3.45286 Test Statistic under Ho X1-A2 7.8-5.2 Paoiea T 2.6 2.4372 345286,--+ 20 22

P-Value: 0.0193

Here P-value < alpha 0.04 so we reject H0

Thus we conclude that The student cards are older than faculty cars

Answer: Yes

b)

Pooled SD ,2 속(M -1) ,2ך119-*3.42+ 21*3.52 20+22-2345286 The 96% confidence interval of the difference of two population means(M-A) is (h-n)±t(004al+t2-an,kdYn+ぬ=(78-52)±2.1229*345286는+22 7.8-5.2) 2. 1229 *3.45286|-+--(0.3353071 , 4.864693) (a3353071 ,4.864693) 20 22